# Parametric Expensive Multi-Objective Optimization via Generative Solution Modeling

### v3

@@ Abstract
Many real-world applications require solving families of expensive multi-objective optimization problems (EMOPs) under varying operational conditions. This gives rise to parametric expensive multi-objective optimization problems (P-EMOPs) where each task parameter defines a distinct optimization instance. Current multi-objective Bayesian optimization methods have been widely used for finding finite sets of Pareto optimal solutions for individual tasks. However, P-EMOPs present a fundamental challenge: the continuous task parameter space can contain infinite distinct problems, each requiring separate expensive evaluations. This demands learning an inverse model that can directly predict optimized solutions for any task-preference query without expensive re-evaluation. This paper introduces a novel parametric multi-task multi-objective Bayesian optimizer that learns this inverse model by alternating between (1) acquisition-driven search leveraging inter-task synergies and (2) generative solution sampling via conditional generative models. This approach enables efficient optimization across related tasks and finally achieves direct solution prediction for unseen parameterized EMOPs without additional expensive evaluations. We theoretically justify the faster convergence by leveraging inter-task synergies through task-aware Gaussian processes. Meanwhile, based on that, empirical studies of our optimizer and inverse model in synthetic and real-world benchmarks further verify the effectiveness of the proposed generative alternating framework.
@@ 1 Introduction
To efficiently identify trade-off solutions, surrogate modeling has become essential for reducing the cost of evaluating expensive objectives. Among these, multi-objective Bayesian optimization (MOBO) stands out as the leading strategy. MOBO methods have been widely used for identifying finite sets of Pareto optimal solutions by scalarizing multiple objectives into single objectives [27, 44, 41] or by optimizing metrics such as expected hypervolume improvement [14, 11, 12] and uncertainty reduction [42, 20, 53, 3]. To move beyond these finite solution‑set approximations and directly model the entire trade‑off surface, recent efforts have extended MOBO to continuous Pareto‑manifold learning [32, 34, 31].
However, in many real-world scenarios, optimization problems arise not as isolated instances but as families of related tasks under varying operational conditions. For example, turbine components must be optimized across different wind speeds [35], and robotic control policies must be optimized for diverse terrains [55] or damage conditions [10]. These cases create fundamental challenges: the continuous range of operational conditions forms infinitely many related optimization problems, each demanding separate, costly evaluations under current MOBO methods. As operational conditions frequently shift [39], repeatedly performing full optimizations to evaluate trade-offs across multiple scenarios becomes prohibitively expensive and computationally infeasible. Therefore, handling families of related EMOPs requires new approaches that avoid excessive evaluation costs. To our best knowledge, little existing work has systematically addressed this parametric setting.
This paper proposes a novel parametric multi-objective optimizer to address this challenge by learning an inverse model. This inverse model can directly predict optimized solutions for any given task parameter, thereby bypassing the repeated costly evaluations mentioned above. Our approach alternates between two complementary phases: (1) acquisition-driven search guided by task-aware Gaussian processes, which explicitly leverages inter-task correlations to efficiently explore the solution space, and (2) generative solution sampling via conditional generative models, which learns high-quality solution distributions conditioned on task parameters and preferences. This dual strategy not only significantly enhances optimization efficiency across multiple related tasks but also finally builds an inverse model for predicting optimized solutions to previously unseen EMOPs without additional expensive evaluations.
@@
  • We design an alternating optimization approach that combines acquisition-driven search with conditional generative solution sampling, enabling both efficient multi-task optimization and inverse model building.
  • We develop a Parametric Multi-Task Multi-Objective Bayesian Optimizer (PMT-MOBO) with task-aware Gaussian Processes, demonstrating its faster convergence over single-task counterpart empirically and theoretically.
@@ 2.2 Parametric Expensive Multi-Objective Optimization
In many real-world scenarios, EMOPs must be solved repeatedly under varying environmental [35] or operational conditions [43]. For example, 3D model design considers optimizing geometric features to maintain the structural performance under varying external physical constraints [45]. These variations can be represented using continuous task parameters $\boldsymbol{\theta}\in\Theta\subseteq\mathbb{R}^{V}$, which defines a family of different EMOPs over the same decision space. This leads to the formulation of a parametric expensive multi-objective optimization problem (P-EMOP):
@@
where each task parameter $\boldsymbol{\theta}$ defines a distinct and expensive multi-objective problem instance $F(\mathbf{x},\boldsymbol{\theta})$. The ideal goal of P-EMOP is to approximate the Pareto Set across potentially infinite task parameters under limited evaluation budgets. We address this by learning an inverse model $\mathcal{M}(\boldsymbol{\theta},\boldsymbol{\lambda})$ that generates optimized solutions for any task parameter $\boldsymbol{\theta}$ and preference vector $\boldsymbol{\lambda}$ without expensive re-optimization. This represents a fundamental shift from per-task optimization to learning generalized solution mappings across the continuous parameter space. Here, $\mathfrak{s}_{\boldsymbol{\lambda}}(\cdot):\mathbb{R}^{M}\rightarrow\mathbb{R}$ is a scalarization function [^1] In this work, we instantiate $\mathfrak{s}_{\boldsymbol{\lambda}}(\cdot)$ for all methods using a recently proposed hypervolume scalarization function [65]. characterized by preference vector $\boldsymbol{\lambda}$. We incorporate $\boldsymbol{\lambda}$ into the inverse model since the Pareto Set generally lies on a continuous manifold for non-trivial multi-objective optimization [37], enabling $\mathcal{M}(\boldsymbol{\theta},\boldsymbol{\lambda})$ to output a single solution vector rather than a solution set.
Although (2) contains infinitely many tasks, these EMOPs share inter-task synergies due to their parametric relationship. Rather than treating tasks in isolation, we solve a finite sample $\{\boldsymbol{\theta}_{1},\dots,\boldsymbol{\theta}_{K}\}$ as a multi-task optimization problem [58], assuming tasks belong to a coherent family sharing solution components or objective landscape regularities [61]. This assumption, pioneered by multi-objective multifactorial optimization [18] and commonly exploited in subsequent multi-task optimization works [19, 2, 57] and recent parametric multi-objective optimization studies [7, 8], enables more efficient solution search across the task parameter space through inter-task synergies, ultimately yielding higher-quality solutions for constructing the inverse model.
@@ 2.3 Gaussian Processes for Multi-Objective Optimization
Algorithm 1 Workflow of ST-MOBO and PMT-MOBO
Output: Final datasets $\{\mathcal{D}_{k}\}_{k=1}^{K}$.
1:for $t=1$ to $T$ do
2:  Sample preference vector $\boldsymbol{\lambda}\sim\Lambda$
3:  for each task $\boldsymbol{\theta}_{k}$ do
4:   if $\mathcal{A}=\texttt{ST-MOBO}$ then
5:    Select $\mathbf{x}_{k}^{(t)}$ using (4) and (6)
6:   else
7:    Select $\mathbf{x}_{k}^{(t)}$ using (5) and (7)
8:   end if
10:  end for
11:  for each objective $f_{m}$ do
12:   if $\mathcal{A}=\texttt{ST-MOBO}$ then
13:    for each task $\boldsymbol{\theta}_{k}$ do
15:     Train $\mathcal{GP}_{m}$ on $\mathcal{D}_{k}^{(m)}$
16:    end for
17:   else
19:    Train $\mathcal{GP}_{TA,m}$ on $\mathcal{D}_{\text{joint}}^{(m)}$
20:   end if
21:  end for
22:end for
23:return $\{\mathcal{D}_{k}\}_{k=1}^{K}$
@@ 3.2 PMT-MOBO with Task-Aware GP
where $\boldsymbol{\widetilde{\alpha}}(\mathbf{x},\boldsymbol{\theta}_{k})\in\mathbb{R}^{M}$ denotes the vector of task-aware acquisition function values for the EMOP task instance parameterized by $\boldsymbol{\theta}_{k}$. We refer to this method as PMT-MOBO. For clarity, a comparative workflow of either ST-MOBO or PMT-MOBO is showcased in Algorithm 1.
@@ 3.3 Theoretical Analysis of Faster Convergence
For our theoretical analysis, we instantiate the acquisition functions $\boldsymbol{\alpha}(\mathbf{x})$ and $\boldsymbol{\widetilde{\alpha}}(\mathbf{x},\boldsymbol{\theta}_{k})$ using the Upper Confidence Bound (UCB) [49]. Specifically, for ST-MOBO, the acquisition function in (4) becomes:
@@
For an expensive multi-objective optimization problem with a total evaluation budget $T$, the goal is to obtain a set of solutions $\mathbf{X}_{T}\subset\mathcal{X}$ that approximates the true Pareto Set $\mathcal{X}^{*}$. Following the analysis technique [41], this goal can be quantified using the notion of Bayes Regret, which evaluates the quality of the obtained solution set with respect to random preference vectors.
@@
In the theoretical study, we assume that each objective function $f_{m}(\cdot)$ follows a Gaussian distribution with marginal variances upper bounded by $1$ and with independent noise variance upper bounded by $\sigma^{2}$ on a compact domain $\mathcal{X}\subseteq[0,1]^{D}$. Moreover, it is also assumed that scalarization functions $\mathfrak{s}_{\boldsymbol{\lambda}}(\cdot)$ is $J_{\boldsymbol{\lambda}}$-Lipschitz ($J_{\boldsymbol{\lambda}}\leq J,\forall\boldsymbol{\lambda}\in\Lambda$, $\Lambda$ being a bounded subset of a normed linear space) in the $\ell_{1}$-norm and monotonically increasing for each objective.
@@
Under Assumption 1 and with the task‐aware kernel that satisfies: $\widetilde{\kappa}_{m}\bigl((\mathbf{x},\boldsymbol{\theta}_{k}),(\mathbf{x}^{\prime},\boldsymbol{\theta}_{k})\bigr)=\kappa_{m}(\mathbf{x},\mathbf{x}^{\prime}),\forall k\in[K],$ let $\widetilde{\gamma}_{T,k}=\max_{m}\widetilde{\gamma}_{T,m,k}$ be the largest MIG of PMT-MOBO across $M$ objective functions over $T$ scalarized evaluations for the $k$-th task, it holds that
$\widetilde{\gamma}_{T,m,k}\leq\gamma_{T,m,k},\forall m\in[M],\forall k\in[K].$ (13)
Accordingly, it also holds that $\widetilde{\gamma}_{T,k}\leq\gamma_{T,k},\forall k\in[K].$ Consider that both regret bounds from Theorem 1 still hold with $\gamma_{T,k}$ replaced by $\widetilde{\gamma}_{T,k}$, consequently, it holds that,
@@
i.e. PMT‐MOBO achieves a tighter upper bound on Bayes regret than standard ST‐MOBO.
The detailed proof of Theorem 2 can be found in the Appendix A.
@@
3:  mode $\leftarrow$ ACQUISITION or GENERATIVE
@@ 3.4 Conditional Generative Model
In this way, distribution‐aware sampling is seamlessly blended with uncertainty‐driven exploration. We instantiate this conditional generative model via either Variational Autoencoder (VAE) [25] or Denoising Diffusion Probabilistic Model (DDPM) [21] as follows.
@@ 3.5 Alternating Optimization Loop
The complete framework alternates between two complementary search modes as illustrated in Figure 1: acquisition-driven search via PMT-MOBO and generative solution sampling using the conditional generative model. Algorithm 2 presents the detailed workflow alternating between ACQUISITION and GENERATIVE modes. We alternate between acquisition-driven and generative sampling modes every iteration in a round-robin fashion.
In ACQUISITION mode, solutions are selected using the task-aware acquisition function. In GENERATIVE mode, candidates are sampled from $p_{\phi}(\mathbf{x}|\boldsymbol{\lambda}^{(t)},\boldsymbol{\theta}_{k})$ and the best is selected using the same acquisition function. Both update task-aware GPs after evaluation, while the generative model gets retrained during GENERATIVE iterations using $\mathcal{D}^{(gen)}$.
@@
We maintain datasets per task-preference pair $(\boldsymbol{\lambda},\boldsymbol{\theta}_{k})$, storing top $Q\%$ solutions by scalarization score $\mathfrak{s}_{\boldsymbol{\lambda}}(F(\mathbf{x},\boldsymbol{\theta}_{k}))$ [40]. Preference vectors $\boldsymbol{\lambda}$ are generated via random sampling from $\Lambda$. Each elite solution $\mathbf{x}$ pairs with conditioning vector $\boldsymbol{c}=(\boldsymbol{\lambda},\boldsymbol{\theta}_{k})$ for generative model training.
Method DTLZ-1 DTLZ-2 DTLZ-3 LAMP SOLAR MAGNETIC UAV
qEHVI 3.23e+04 (5.06e+01)+- 2.33e+00 (6.20e-02) 2.66e+04 (1.85e-02)+- 6.14e-01 (3.00e-03) 6.07e-01 (1.29e-02) 2.18e-01 (1.15e-02)++ 4.89e-01 (2.60e-03)
PPSL-MOBO 3.20e+04 (1.72e+03)+- 2.86e+00 (3.00e-01)≈- 2.71e+04 (7.89e+02)+- 6.09e-01 (1.61e-02) 6.35e-01 (1.15e-02) 9.41e-02 (1.69e-02) 5.85e-01 (3.19e-02)≈-
PMT-MOBO 1.71e+04 (4.70e+03) 2.77e+00 (5.50e-02) 3.78e+03 (5.40e+03) 5.83e-01 (1.22e-02) 7.10e-01 (1.90e-02) 1.73e-01 (2.99e-02)+- 4.78e-01 (1.27e-02)
Table 1: Average Hypervolume results (mean $\pm$ std) across synthetic and real-world benchmarks. Statistical significance is assessed using the Wilcoxon signed-rank test at the 90% confidence level. Superscripts $^{s_{1}s_{2}}$ indicate that the competitor is significantly worse (-), significantly better (++), or statistically similar ($\approx$) relative to PMT-MOBO-VAE ($s_{1}$) and PMT-MOBO-DDPM ($s_{2}$).

### v4

@@ Abstract
Many real-world applications require solving families of expensive multi-objective optimization problems (EMOPs) under varying operational conditions. This can be formulated as parametric expensive multi-objective optimization problems (P-EMOPs) where each task parameter defines a distinct optimization instance. Current multi-objective Bayesian optimization methods have been widely used for finding finite sets of Pareto optimal solutions for individual tasks. However, P-EMOPs present a fundamental challenge: the continuous task parameter space can contain infinite distinct problems, each requiring separate expensive evaluations. This demands learning an inverse model that can directly predict optimized solutions for any task-preference query without expensive re-evaluation. This paper introduces a novel parametric multi-objective Bayesian optimizer that learns this inverse model by alternating between (1) generative solution sampling via conditional generative models and (2) acquisition-driven search leveraging inter-task synergies. This approach enables effective optimization across multiple tasks and finally achieves direct solution prediction for unseen parameterized EMOPs without additional expensive evaluations. We theoretically justify the faster convergence by leveraging inter-task synergies through task-aware Gaussian processes. Based on that, empirical studies in synthetic and real-world benchmarks further verify the effectiveness of the proposed parametric optimizer.
@@ 1 Introduction
To efficiently identify trade-off solutions, surrogate modeling has become essential for reducing the cost of evaluating expensive objectives. Among these, multi-objective Bayesian optimization (MOBO) stands out as the leading strategy. MOBO methods have been widely used for identifying finite sets of Pareto optimal solutions by scalarizing multiple objectives into single objectives [26, 39] or by optimizing metrics such as expected hypervolume improvement [13, 11] and uncertainty reduction [40, 19, 49, 3]. To move beyond these finite solution‑set approximations and directly model the entire trade‑off surface, recent efforts have extended MOBO to continuous Pareto‑manifold learning [30, 29, 8].
However, in many real-world scenarios, optimization problems arise not as isolated instances but as families of related tasks under varying operational conditions. For example, turbine components must be optimized across different wind speeds [33], and robotic control policies must be optimized for diverse damage conditions [10]. These cases create fundamental challenges: the continuous range of operational conditions forms infinitely many related optimization problems, each demanding separate, costly evaluations under current MOBO methods. As operational conditions frequently shift [37], repeatedly performing full optimizations to evaluate trade-offs across multiple scenarios becomes prohibitively expensive and computationally infeasible. Therefore, handling families of related EMOPs requires new approaches that avoid excessive evaluation costs. To our best knowledge, little existing work has systematically addressed this parametric setting.
This paper proposes a novel parametric multi-objective Bayesian optimizer to address this challenge by learning an inverse model. This inverse model can directly predict optimized solutions for any given task parameter, thereby bypassing the repeated costly evaluations mentioned above. Our approach alternates between two complementary phases: (1) acquisition-driven search guided by task-aware Gaussian processes, which explicitly leverages inter-task correlations to efficiently explore the solution space, and (2) generative solution sampling via conditional generative models, which learns high-quality solution distributions conditioned on task parameters and preferences. This dual strategy not only significantly enhances optimization efficiency across multiple related tasks but also finally builds an inverse model for predicting optimized solutions to previously unseen EMOPs without additional expensive evaluations.
@@
  • We design an alternating optimization approach that curates the acquisition-driven search with conditional generative solution sampling, enabling both efficient multi-task optimization and inverse model building.
  • We develop a Parametric Multi-Task Multi-Objective Bayesian Optimizer (PMT-MOBO) with task-aware Gaussian Processes, demonstrating its faster convergence over single-task counterparts empirically and theoretically.
@@ 2.2 Parametric Expensive Multi-Objective Optimization
In many real-world scenarios, EMOPs must be solved repeatedly under varying environmental [33] or operational conditions [41]. These variations can be represented using continuous task parameters $\boldsymbol{\theta}\in\Theta\subseteq\mathbb{R}^{V}$, which defines a family of different EMOPs over the same decision space. This leads to the formulation of a parametric expensive multi-objective optimization problem (P-EMOP):
@@
where each task parameter $\boldsymbol{\theta}$ defines a distinct and expensive multi-objective problem instance $F(\mathbf{x},\boldsymbol{\theta})$. The ideal goal of P-EMOP is to approximate the Pareto Set across potentially infinite task parameters under limited evaluation budgets. We address this by learning an inverse model $\mathcal{M}(\boldsymbol{\theta},\boldsymbol{\lambda})$ that generates optimized solutions for any task parameter $\boldsymbol{\theta}$ and preference vector $\boldsymbol{\lambda}$ without expensive re-optimization. This represents a fundamental shift from per-task optimization to learning generalized solution mappings across the continuous parameter space. Here, $\mathfrak{s}_{\boldsymbol{\lambda}}(\cdot):\mathbb{R}^{M}\rightarrow\mathbb{R}$ is a scalarization function characterized by preference vector $\boldsymbol{\lambda}$. We incorporate $\boldsymbol{\lambda}$ into the inverse model since the Pareto Set generally lies on a continuous manifold for non-trivial multi-objective optimization [35], enabling $\mathcal{M}(\boldsymbol{\theta},\boldsymbol{\lambda})$ to output a single solution vector rather than a solution set.
Although (2) contains infinitely many tasks, these EMOPs share inter-task synergies due to their parametric relationship. Rather than treating tasks in isolation, we solve a finite sample $\{\boldsymbol{\theta}_{1},\dots,\boldsymbol{\theta}_{K}\}$ as a multi-task optimization problem [53], assuming tasks belong to a coherent family sharing solution components or objective landscape regularities. This assumption, pioneered by multi-objective multifactorial optimization [17] and commonly exploited in subsequent multi-task optimization works [18, 52, 2] and recent parametric multi-objective optimization studies [7], enables more efficient solution search across the task parameter space through inter-task synergies, thereby yielding higher-quality solutions for constructing the inverse model.
@@ 3.2 PMT-MOBO with Task-Aware GP
where $\boldsymbol{\widetilde{\alpha}}(\mathbf{x},\boldsymbol{\theta}_{k})\in\mathbb{R}^{M}$ denotes the vector of task-aware acquisition function values for the EMOP task instance parameterized by $\boldsymbol{\theta}_{k}$. We refer to this method as PMT-MOBO. For clarity, a comparative workflow of either ST-MOBO or PMT-MOBO is showcased in Algorithm 2 of Appendix F.
@@ 3.3 Theoretical Analysis of Faster Convergence
For our theoretical analysis, we instantiate the acquisition functions $\boldsymbol{\alpha}(\mathbf{x})$ and $\boldsymbol{\widetilde{\alpha}}(\mathbf{x},\boldsymbol{\theta}_{k})$ using the Lower Confidence Bound (LCB) [45]. These LCB instantiations coincide with the acquisition rules employed in our practical ST-MOBO and PMT-MOBO algorithms, so the following regret analysis directly characterizes the convergence behavior of the implemented methods. Specifically, for ST-MOBO, the acquisition function in (4) becomes:
@@
For an EMOP with a total evaluation budget $T$, the goal is to obtain a set of solutions $\mathbf{X}_{T}\subset\mathcal{X}$ that approximates the true Pareto Set $\mathcal{X}^{*}$. Following the analysis technique [39], this goal can be quantified using the notion of Bayes Regret, which evaluates the quality of the obtained solution set with respect to random preference vectors.
@@
For each objective $f_{m}$ and each task parameter $\boldsymbol{\theta}_{k}$, the function $\mathbf{x}\mapsto f_{m}(\mathbf{x},\boldsymbol{\theta}_{k})$ satisfies the same regularity and GP assumptions as in Theorem 1 of [39] (independent sub-Gaussian noise on a compact domain, and scalarizations that are monotone and Lipschitz in all objectives).
@@
Under Assumption 1, consider a PMT-MOBO model where the task-aware kernel yields covariance values equivalent to the single-task kernel for any pair of inputs belonging to the same task. Let $\widetilde{\gamma}_{T,k}=\max_{m\in[M]}\widetilde{\gamma}_{T,m,k}$ be the largest MIG of PMT-MOBO across $M$ objectives over $T$ scalarized evaluations for task $k$.
Then, for all $m\in[M]$ and $k\in[K]$, it holds that
$\widetilde{\gamma}_{T,m,k}\;\leq\;\gamma_{T,m,k},$ (13)
and accordingly $\widetilde{\gamma}_{T,k}\leq\gamma_{T,k}$ for every $k\in[K]$. Since the proof of Theorem 1 depends on each task only through its MIG term, the same regret bounds remain valid with $\gamma_{T,k}$ replaced by $\widetilde{\gamma}_{T,k}$, and consequently
@@
i.e., PMT-MOBO achieves a tighter upper bound on Bayes regret than standard ST-MOBO.
Theorem 2 establishes that PMT-MOBO achieves a tighter Bayes regret bound than ST-MOBO. The detailed proof (Appendix A) demonstrates that conditioning on observations from related tasks strictly reduces the posterior covariance for the target task. Crucially, under the same regularity assumptions utilized in [39], one can ensure that the structural form of the original regret upper bound remains valid. Since this bound is a monotonically increasing function of the Maximum Information Gain (MIG), the inequality, $\widetilde{\gamma}_{T,k}\leq\gamma_{T,k}$ derived from the reduced posterior variance, translates to a strictly lower upper bound on the cumulative regret.
@@
3:  mode $\leftarrow$ ACQUISITION or GENERATIVE (Round-Robin)
@@ 3.4 Conditional Generative Model
In this way, distribution‐aware sampling is seamlessly blended with uncertainty‐driven exploration. We instantiate this generative model via either Variational Autoencoder (VAE) [24] or Denoising Diffusion Probabilistic Model (DDPM) [20] as follows.
@@ 3.5 Alternating Optimization Loop
The complete framework alternates between two complementary search modes as illustrated in Figure 1: acquisition-driven search via PMT-MOBO and generative solution sampling using the conditional generative model. Algorithm 1 presents the detailed workflow alternating between ACQUISITION and GENERATIVE modes. We alternate between generative sampling and acquisition-driven modes every iteration in a round-robin fashion. In GENERATIVE mode, candidates are sampled from $p_{\phi}(\mathbf{x}|\boldsymbol{\lambda}^{(t)},\boldsymbol{\theta}_{k})$ and the best is selected using the acquisition function. In ACQUISITION mode, solutions are directly selected by optimizing the task-aware acquisition function. Both update task-aware GPs after evaluation, while the generative model gets retrained during GENERATIVE iterations using $\mathcal{D}^{(gen)}$.
@@
We maintain datasets per task-preference pair $(\boldsymbol{\lambda},\boldsymbol{\theta}_{k})$, storing top $Q\%$ solutions by scalarization score $\mathfrak{s}_{\boldsymbol{\lambda}}(F(\mathbf{x},\boldsymbol{\theta}_{k}))$ [38]. Preference vectors $\boldsymbol{\lambda}$ are generated via random sampling from $\Lambda$. Each elite solution $\mathbf{x}$ pairs with conditioning vector $\boldsymbol{c}=(\boldsymbol{\lambda},\boldsymbol{\theta}_{k})$ for generative model training. More detailed settings can be found in Appendix C.3.
Method DTLZ-1 DTLZ-2 DTLZ-3 LAMP SOLAR UAV
PSL-MOBO 2.82e+04 (1.78e+03)≈- 1.90e+00 (1.30e-01) 2.39e+04 (3.22e+03)+- 5.01e-01 (7.32e-02) 3.54e-01 (4.06e-02) 4.12e-01 (7.30e-03)
CDM-PSL 1.92e+04 (3.85e+03) 2.52e+00 (8.49e-02) 1.83e+03 (1.31e+03) 6.19e-01 (4.45e-02)≈- 6.54e-01 (3.11e-02) 4.26e-01 (3.09e-02)
PPSL-MOBO 2.24e+04 (1.10e+02) 2.98e+00 (3.56e-02) 2.67e+04 (6.08e+02)+- 6.12e-01 (1.05e-01)≈- 4.86e-01 (1.11e-01) 5.24e-01 (1.71e-01)
PMT-MOBO-VAE 2.88e+04 (2.57e+03) 3.10e+00 (2.99e-02) 1.78e+04 (7.28e+03) 6.31e-01 (7.33e-02) 7.93e-01 (1.56e-02) 6.08e-01 (1.88e-02)
PMT-MOBO-DDPM 3.47e+04 (4.85e+02) 3.13e+00 (1.41e-02) 3.20e+04 (1.98e+03) 6.70e-01 (5.82e-02) 7.65e-01 (1.30e-02) 6.05e-01 (2.44e-02)
Table 1: Inverse Model Generalization Performance on Unseen EMOP Tasks. Statistical significance is assessed using the Wilcoxon signed-rank test at the 95% confidence level. Superscripts $^{s_{1}s_{2}}$ indicate that the competitor is significantly worse (-), significantly better (++), or statistically similar ($\approx$) relative to PMT-MOBO-VAE ($s_{1}$) and PMT-MOBO-DDPM ($s_{2}$).